OpenStudy (lucaz):

derivative of y = (x^2 + 1)e^x ?

OpenStudy (lucaz):

I found this answer e^x(x^2+1) + e^x2x

OpenStudy (raden):

u are right

OpenStudy (lucaz):

my says (x+1)^2e^x

OpenStudy (lucaz):

my book..

OpenStudy (raden):

that's same like yours

OpenStudy (raden):

it just simplify form by completing quadratic equation: e^x(x^2+1) + e^x2x = e^x ( x^2+1+2x) = e^x (x+1)^2

OpenStudy (lucaz):

yeah, now I see, thanks

OpenStudy (raden):

you're welcome